15: Solve the following equations.
12m2 + 7m = -1

Answer by tutorcecilia(2152)

12m2 + 7m = -1 [Add (1) to both sides]
12m^2+7m+1=0
(4m+1)(3m+1)=0
.
4m+1=0 [set each factor equal to zero and solve for the x-term]
4m=-1
m=-1/4
and
3m+1=0
3m=-1
m=-1/3
.
check by plugging either value of x back into the original equation and solve.